A particular brand of diet margarine has analysed to determine the level of polyunsaturated fatry acids (in percentages).A sample of six packages resulted in the following data 16.8,17.2,17.2,16.9,16.5,17.1 .Find the 95%confidence interval for the mean level of polyunsaturated fatty accid for this brand of diet margarine


1
Expert's answer
2021-05-19T16:41:38-0400

mean "(\bar{x})=\dfrac{\sum x}{n}=\dfrac{1019}{6}=16.9833"


Standard deviation "(S)=\sqrt{\dfrac{1}{(n-1)}(\sum X^2-\dfrac{(\sum x)^2}{n})}"


           "= \sqrt{\dfrac{1}{5}(11731.11-\dfrac{(101.91)^2}{6})}=0.3189"



at "\alpha=0.05, z_{\frac{\alpha}{2}}=z_{0.05}=1.96"


So 95% confidence interval is "=\bar{x}\pm z_{0.025}\times \dfrac{S}{\sqrt{n}}"

               "=16.9833\pm(1.96\times \dfrac{0.3189}{\sqrt{6}}
\\[9pt]
 =(16.9833\pm 0.255)
\\
 =(16.728, 17.2383)"